a) An aqueous reactant stream (4mol/liter) passes through a mixed flow reactor followed by a plug flow reactor. Find the concentration at the exit of the plug flow reactor if in the mixed flow reactor CA-Imol/liter. The reaction is second onder with respect to A: the volume of the plug flow reactor is three times that of the mixed flow reactor. What is the effect on the exit concentration when the reactor combination is neversed for the same feed rate and by how much?


1
Expert's answer
2021-08-31T01:42:40-0400

For the second-order reaction

for the mixed flow reaction

"R_m = \frac{C_A0-C_A1}{C_A1^2}\\
R_m = \frac{4-1}{1}=3\\"

So for the plug flow reactor

"RT_p= 3 kT_m = 9\\
\frac{C_A2}{C_A1}= \frac{1}{1+kT_pC_A1}\\
\frac{C_A2}{C_A1}=\frac{1}{1+9(1)}\\
\frac{C_A2}{C_A1}=0.1\\
C_A2=0.1C_A1=0.1 mol/lit"


The reverse for the same feed rate has no effect.